OpenStudy (anonymous):

pi= Delta when pi=4 deltaR =10

OpenStudy (anonymous):

since pi = delta then you can write it like this piR = 10 since pi=4 4R = 10 so R = 10/4 = 5/2

OpenStudy (anonymous):

thanks hun

OpenStudy (anonymous):

np glad to help

OpenStudy (anonymous):

u can help me more because nobody else on here helping me

OpenStudy (anonymous):

yea post em up

OpenStudy (anonymous):

rd

OpenStudy (anonymous):

whats the value of x must be inserted into the following wquation so that y has the value of 5 y=2x

OpenStudy (anonymous):

so plug in the y value and solve for x 5 = 2x x = 5/2

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

yea x is 5/2

OpenStudy (anonymous):

thanks here is another one what must x be so that the dirrerence between y and z is 2.5 y-z=10x

OpenStudy (anonymous):

you want y-z = 2.5 so replace that and get 2.5 = 10x solve for x x = 0.25

OpenStudy (anonymous):

thanks for yur help

OpenStudy (anonymous):

np glad to help

OpenStudy (anonymous):

:)